• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Parno

    Log entry time 12:49:39 on December14,2008

    Entry number 253383

    This entry is a followup to: 253354

    Followups:

    keyword=Compton rates were high because of slit adjustments

    The high Compton rates last night, and resulting loss of Compton spectrum / asymmetry, were due to slit adjustments at MCC. See ELOG entry 1449456. Slits were readjusted from -1 to 29.6 at around 11:40 last night, solving the rate problem and bringing the Compton spectrum back. Mystery solved!